OpenStudy (anonymous):

A bike shop sells rolling strollers for runners with babies, with a total of 28 in the store. Some of the strollers have 3 wheels and some have 4. If there are 100 wheels, how many 3-wheel strollers are there? A. 25 B. 12 C. 3 D. 16

OpenStudy (anonymous):

let x = number of 3 wheel stollers let y = number of 4 wheel strollers x + y = 28 y = 28 - x 3x + 4y = 100 3x + 4(28 - x) = 100 3x + 112 - 4x = 100 -x = -12 x = 12 B) there are 12, three wheel strollers